OpenStudy (anonymous):

What is the simplified form of (x^2-25)/(x-5) x – 5, with the restriction x ≠ –5 x + 5, with the restriction x ≠ 5 x + 5, with the restriction x ≠ –5 x – 5, with the restriction x ≠ 5

OpenStudy (help!!!!):

i got x+5

OpenStudy (help!!!!):

x – 5, with the restriction x ≠ 5

OpenStudy (anonymous):

thanks :)

OpenStudy (help!!!!):

medal?

OpenStudy (ghazi):

\[\frac{ x^2-25 }{ x-5 }=\frac{ (x+5)(x-5) }{ x-5 }=(x+5)\] so its x+5 with x not equal to 5 ( B) else it will be infinity